OpenStudy (anonymous):

the vertex of a function y=9-8x-x^2 is (-4,25)?

OpenStudy (anonymous):

The vertex can be found by the x co-ordinate being, \[\frac{ -b }{ 2a }\] than just substitute back to find the y co-ordinate.

OpenStudy (ybarrap):

watch the sign

OpenStudy (ybarrap):

remember, a = -1

OpenStudy (anonymous):

@some_someone i cant message you unless you fann me :P

OpenStudy (anonymous):

hello again

OpenStudy (ybarrap):

\(y=- (b^{2}-4ac)/4a\)

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

guys i know you guys are trying to help but like ive never done this so idk what you guys are doing .. :(

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

it is in fact \((-4,25)\)

OpenStudy (agent0smith):

y=9-8x-x^2 first rewrite this as y= ax^2 +bx + c: y = -x^2 - 8x + 9 now compare to y= ax^2 +bx + c

OpenStudy (agent0smith):

So a=-1. b=-8. Now use -b/(2a) to find the x-coordinate of the vertex.

OpenStudy (ybarrap):

The equation you have is a quadratic. It is a "bowl." Because of the negative sign in front of the quadratic term, the bowl is "upside down" opening up to the bottom. So there is a peak somewhere above this bowl. This peak occurs at (-4,25). Try any point to the left and right of x=-4 and y will always be smaller. That's how you know you found the peak, or the vertex of this parabola.